OpenStudy (anonymous):

Express the number in a/b form where a and b are integers. (-.125)^2/3

OpenStudy (luigi0210):

\[\LARGE -\frac{1}{8}=-0.125\]

OpenStudy (anonymous):

what

OpenStudy (triciaal):

is that (-0.125)^2/3? to change from decimal to fraction put a 1 under the decimal point and add as many 0 as significant figures after in the case -0.125 = -125/1000 this can be simplified factors of 125 =5x5x5 factors of 1000 2x2x2x5x5x5 so divide by (cancel) 5x5x5 to get -1/8 now you have (-1/8)^2/3 from the rules of exponents same as (-1/8)^1/3 then ^2 8 can be written as 2^3 as above (-1/2^3)^1/3 = (-1/2)^1 =-1/2 (-1/2)^2 = 1/4 negative times negative is positive note that if you recognize the factors you can always do the powers first (125/1000)^2/3 125 = 5^3 1000 = 10^3 (-5/10)^3 raised to the 2/3 =(-1/2)^(3)(2/3) =(-1/2)^2 =1/4 my answer is 1/4